# More than 200 Globular Clusters in the Milky Way, and still no super-Solar metallicity ones

### v1

@@ Abstract
Context. Many globular clusters in the Milky Way have been studied in recent years, especially in hidden regions such as those of the Galactic bulge.
Aims. The main goal of this paper is to understand what we can learn if we include these new objects into the Milky Way Globular Cluster system that we know today. We catalogue 37 recently discovered globular clusters. Most of them are located in the Galactic bulge, but we also include some disk’s globular clusters for comparison.
Methods. We use different distributions for investigating the Galactic globular cluster system: the metallicity distribution, the luminosity function, and the age distribution. We consider three samples. We first treat separately the new globular clusters sample from the known and well-characterised globular clusters. Consequently, we merge these two samples, thus upgrading the Milky Way globular cluster system. Furthermore, we performed a comparison between our clusters sample and the field star population.
Results. We find a double peaked distribution for the luminosity function, which shows an elongated faint end tail. Considering the ”merged” sample, the luminosity function peaks at $M_{V}^{up}=-7.00\pm 1.3$ mag and at $M_{V}^{up}=-4.1\pm 0.48$ mag. The metallicity distributions also display a bimodality trend. In this case, we compare our new sample compilation with previously published ones, finding that the distributions are in good general agreement. We also construct the metallicity distribution for the field star sample, and comparing this with that one of the globular clusters, we learn that a high percentage of field stars show [Fe/H] $>0$, whereas we do not detect any globular clusters in the same metallicity range. In order to understand this inconsistency, we construct the age-metallicity diagram for both samples, noting that the old and metal-poor population (age $\geq 8$ Gyr and [Fe/H] $\leq-1.0$) is represented by globular clusters, while the young and metal-rich population (age $<8$ Gyr and [Fe/H] $>-1.0$) corresponds to field stars.
Conclusions. From the analysis of the globular cluster luminosity function and metallicity distribution, we can conclude that many GCs, probably those very faint, have survived strong dynamical processes, typical of the Bulge regions. Moreover, we cannot exclude the possibility that some of them have been accreted during past merging events, especially the metal-poor component, whereas the metal-rich population may be related to the formation of the bulge and/or disk. Finally, the difference that we notice between the cluster and field star samples should be sought in the evolutionary difference between these two stellar populations.
@@ 1 Introduction
In a cosmological context, the hierarchical galaxy formation process predicts that galaxies and their haloes have a large number of progenitors at the formation redshift ($z\sim 3$) of globular clusters (GCs, De Lucia & Blaizot 2007; Lamers et al. 2017). With the advent of James Webb Space Telescope observations, new theories may be formulated, and new observational discoveries can be done, which may change, confirm or improve our knowledge. For instance, recent works discovered (mini-)quenched galaxies at redshift $z=4.4$ (Looser et al. 2023a) and $z=7.3$ (Looser et al. 2023b), so we cannot exclude the fact that GCs have been formed as well at these redshifts. Also, Mowla et al. (2022) found compact red companions to the Sparkler galaxy, which may be the first evolved GCs seen at $z_{spec}\approx 1.4$. They infer formation redshift of $z_{form}\sim 7-11$ for these candidates, corresponding to ages of $\sim 3.9-4.1$ Gyr at the epoch of observation and a formation time just $\sim 0.5$ Gyr after the Big Bang. Therefore, as they wrote, if these ages are confirmed, at least some GCs appear to have formed contemporaneously with the large-scale reionization of the intergalactic medium, hinting at a deep connection between GCs formation and the initial phases of galaxy assembly. Nevertheless, a long way must be done to at least confirm the existence of those primordial GCs and constructing new models that describe these compact objects. Especially because the exact formation redshifts are somewhat uncertain, they are unresolved and red objects, and no spectroscopic data are still available in order to hint the real phase of these objects.
@@
Both simulations (e.g., Tonini 2013; Kruijssen 2015) and observations (e.g., Horta et al. 2023) showed that much of the metal-poor GC population in massive galaxies was mostly accreted from cannibalised dwarf galaxies, whereas the metal-rich population formed preferentially in-situ [^1] We use ”in-situ” to refer to the GCs that formed during the initial collapse of the central galaxy, and ”ex-situ” for those one that are accreted during the subsequently mergers. or likely during merging events (Hughes et al. 2019a; Garro et al. 2021a; Hasselquist et al. 2021). In this context, debris from accreted GCs and dwarf galaxies are expected to populate halo regions as well as the central regions of the MW (Das et al. 2020; Fernández-Trincado et al. 2019, 2021a). However, the number of confirmed merging footprints in the inner regions is still smaller than those found in the Galactic halo. For that reason, we want to investigate the Galactic bulge in order to collect clues that can help to figure out its formation and evolution history.
@@
Actually, there are many queries still unanswered about the bulge. For instance, it is not completely clear if the bulge is a simple classical bulge, formed from a monolithic collapse, and then grown up hierarchically by merging of galaxies (Guedes et al. 2013), or it was formed by secular evolution (Kormendy & Kennicutt 2004), so it is part of the peanut structure corresponding to the old Galactic thick disk. This would imply that the MW is a pure thin+thick disk galaxy, with only a possible limited contribution by a classical bulge (Di Matteo et al. 2015). Such a mix would represent a compound bulge (Athanassoula 2005; Erwin et al. 2014; Zoccali et al. 2014). Secondly, there are also queries about its evolution, such as how many accretions occurred, how the star formation history and so the chemical evolution proceeded, and so on.
@@
The main aim of this paper is to take advantage from bulge GC populations, which can be used as archaeological probes to reconstruct the formation and evolution of the MW. However, regions like the Galactic bulge are very challenging to study, because of the high stellar crowding and differential reddening, which make the identification of new GCs a complicated task. Wide-field near-infrared (NIR) photometry such as the VISTA Variables in the Vía Láctea (VVV; Minniti et al. 2010; Saito et al. 2012) survey and their extension (VVVX; Minniti 2018) offers a huge advantage to detect several sparse and faint GCs heavily obscured by interstellar dust in the bulge region. Indeed, numerous ($\sim 300$) GC candidates have recently been discovered in these zones (Minniti et al. 2017, 2021c, 2021a; Dias et al. 2022; Garro et al. 2021b, 2022c). It seems that the number of known MW GCs (in total $N_{GC}\sim 170$ Vasiliev & Baumgardt 2021), but in particular the MW bulge GCs ($N_{GC}\sim 40$ Barbuy et al. 2018b, or maybe more as suggested by Bica et al. 2016) is too small if we take into account the mass of the MW and if we compare that number with similar systems, such as M31 galaxy. In this context, for example Abdullah & Kroupa (2018) (see their Figure 6) showed that the MW and M31 specific frequencies should be similar in their range of masses ($M/M_{\odot}=10^{10}-10^{11}$). Therefore, this may imply that the census of the MW clusters is still incomplete (Minniti et al. 2017). Indeed, searching and analysing new GCs in the MW bulge is tricky. They appear as round concentrations of stars, showing up as overdensities above the background in the optical and NIR images, but sometimes they are not real clusters but asterisms, background fluctuations, or they may be streams of previously disrupted GCs. At the distance of the Galactic bulge, their typical half-light radii would range between 2 pc ($800^{\prime}.$80) for the smallest ones and 10 pc ($204^{\prime}.$20) for the largest ones. Additionally, since many of them are very faint and small objects, they are experiencing destruction processes, due to the fact that the deep potential well of the Galactic bulge strengthens the efficiency of such dynamic processes (Habibi et al. 2014; Khoperskov et al. 2018). Hence, this is why the Galactic bulge is referred to as the elephant graveyard (Minniti et al. 2017).
@@
Hence, we want to employ both known and well characterised GCs and the recently discovered ones for investigating the inner regions, using different tools: the metallicity distribution function (MD), the luminosity function (LF), age distributions, and the kinematics information. Additionally, we prefer expanding this analysis to the bulge field stars (FS) populations, since it is expected that a small fraction of bulge FS (proposed to be $\sim 5-10\%$Kruijssen 2012; Adamo et al. 2015) is born in bound systems, and then it is lost due to both merging events and/or destruction/dissolution processes (stellar evolution, two-body relaxation, tidal evaporation, and tidal shocks – Kruijssen et al. 2011; Madrid et al. 2017). Consequently, a comparison between these two populations may help us to understand and constrains the formation and evolution of the inner Galaxy. Nevertheless, these two populations are usually treated separately.
@@
What has been done so far? Different studies have reported that the Galactic bulge shows multiple peaks in its MD, features that are usually associated with different stellar populations (Ness et al. 2013; Bensby et al. 2017; Rojas-Arriagada et al. 2020), coexisting in the innermost regions. The existence of a metallicity gradient is demonstrated adopting different tracers (Minniti et al. 1995). Gonzalez et al. (2013) using the VVV photometry confirmed that metal-rich stars ([Fe/H] $\sim 0$) dominated the inner bulge in regions closer to the Galactic plane ($|b|<5$), while at larger scale heights, the mean metallicity of the bulge red giant population becomes significantly more metal-poor. Ness et al. (2013) and Debattista et al. (2017) found that the metal-rich population is dominant at low latitudes, but drops off rapidly with increasingly latitude. The metal-poor population instead follows the opposite trend, increasing toward high Galactic latitudes. The intermediate-metallicity population dominates everywhere except near the disk, where the metal-rich population dominates. On the other hand, works by Barbuy et al. (2018a) and Savino et al. (2020) found a predominance of moderately metal-poor populations in the innermost Galactocentric regions, which extends to large distances above the Galactic plane. This is probably related to the fact that the old and intermediate age RR Lyrae populations were taken into account, and they should not exhibit a strong peanut shape, as red clump stars (Gonzalez et al. 2015), since they should be older than the formation of the bar itself (Debattista et al. 2017). On the other side, if we take into account GCs, we find the metal-rich ones especially in the inner regions of the Galaxy, while the metal-poor one in the halo.
@@
As we can note, the meaning of ”metal-poor” and ”metal-rich” is not completely defined, and depends on the different contexts where it is used. For instance, for the ”metal-rich” FS we often refer to solar metallicity ([Fe/H] $\approx 0$), while for the ”metal-rich” GCs, we consider values of [Fe/H] $>-0.5$, which might be referred to as ”intermediate-metallicity” for the field. Therefore, in the present work, we broadly define ”metal poor” if [Fe/H] $<-1.0$, whereas ”metal-rich” if [Fe/H] $\geq-1.0$, both for FS and GCs. Moreover, we refer to ”super metal-rich” or ”super-Solar” to indicate those populations with [Fe/H] $\geq 0$. Various nomenclatures are often used in the literature, such as ”moderately metal-poor” (which may correspond to $-2.0\lesssim$ [Fe/H] $\lesssim-1.0$) or ”intermediate metallicity” (which may be equivalent to $-1.5\lesssim$ [Fe/H] $\lesssim-0.6$), but once again it depends on the populations involved. We prefer not to use this nomenclature.
@@
Today, we know of $\sim 170$ well-characterised GCs in the MW (Vasiliev & Baumgardt 2021). Broadly, the Galactic GC system shows a bimodal metallicity distribution, which indicates that in the MW coexist two major subpopulations, but always the metallicity is typically sub-Solar ([Fe/H] $<0$). In spite of that, there are some GCs that may show [Fe/H] $>0$. This is not completely demonstrated, and the misgivings are strongly related to the abundance scale’s calibrations adopted. Therefore, we cannot ignore them. For example, Carretta et al. (2009) suggested a [Fe/H] $>+0.40\pm 0.09$ for Liller 1 GC. Recently, Dalessandro et al. (2022) and Crociati et al. (2023) showed a bimodal iron distribution, with a sub-solar component ([Fe/H] $=-0.48$ dex with $\sigma=0.22$, and age $\approx 12$ Gyr) and a super-solar component ([Fe/H] $=+0.26$ dex with $\sigma=0.17$, and age $\approx 1-3$ Gyr) for Liller 1. For that, they suggested that Liller 1 experienced three main bursts and a low, but constant, activity of star formation over the entire lifetime. Also, NGC 6528 may have a super-solar metallicity as found by Carretta et al. (2009) ([Fe/H] $=+0.07\pm 0.08$) and Liu et al. (2017) ([Fe/H] $=+0.04\pm 0.07$), but in both cases, the uncertainties are larger than the value itself, therefore we cannot conclude that we may or may not have a case of super metal-rich GC.
It is clear that the analysis of the Bulge MD is very complex, since this strongly depends on target selection, tracers and analysis methods.
@@
Moreover, the $\alpha$-elements are efficient tracers of Galactic evolution, especially when [$\alpha$/Fe]-[Fe/H] diagrams are used. An $\alpha$-enhanced plateau at low [Fe/H] is typical both for the MW and external galaxies, this is due to the fast enrichment from massive ($408<M/M_{\odot}<$ 40) stars that explode as type II supernovae (SNII), while a following decline is expected once low-mass (in binary systems with $81.4<M/M_{\odot}<$ 8) stars begin to explode as type Ia supernovae (SNIa). This creates the ”knee” point, which can move to higher metallicity for higher star formation rates (Matteucci & Greggio 1986). For example, Queiroz et al. (2021) found an apparent chemical bimodality in key abundance ratios, such as [$\alpha$/Fe], [C/N], and [Mn/O], which probe different enrichment timescales, indicating a star formation gap between the $\alpha$-rich and $\alpha$-poor populations. They also found signatures of the bar, the thin inner disk, an inner thick disk, and a broad metallicity population with large velocity dispersion indicative of a pressure-supported component. Furthermore, they found a group of counter-rotating stars that could be an accreted metal-poor population originated during a gas-rich accretion phase in the early formation of the bulge, or of a clumpy star formation during the earliest phases of the early disk that migrated into the bulge (Fiteni et al. 2021). In summary, many observations suggest that the formation of the boxy bulge occurred later, originating from the vertical instability of the Galactic bulge, formed via secular evolution from the stellar disk. In this scenario, the old and metal-poor stellar population dominates at higher scale heights due to the non-mixed orbits of originally hotter thick disk stars. Lian et al. (2020) derived three-phase star formation history of the MW bulge, which consists of an initial starburst, followed by a rapid star formation quenching episode and a lengthy, quiescent secular evolution phase. In fact, they found that the metal-poor, high-$\alpha$ bulge stars ([Fe/H] $<0.0$ and [Mg/Fe] $>0.15$) were formed rapidly ($<2$ Gyr) during the early starburst. The density gap between the high- and low-$\alpha$ sequences is due to the quenching process. The metal-rich, low-$\alpha$ population ([Fe/H] $>0.0$ and [Mg/Fe] $<0.15$) then accumulates gradually through inefficient star formation during the secular phase. This is qualitatively consistent with the early star formation history (SFH) of the inner disk. Additionally, this explains why there is a fraction of young stars with age $<5$ Gyr in the bulge. Indeed, we know that the bulge mean age is about 10 Gyr, with a lower-limit of $\sim 5$ Gyr (Clarkson et al. 2011), where well-mixed clusters and stellar populations coexist. The typical age of GCs in the innermost regions is older than 10 Gyr (Harris 1996; Minniti 1995), however GCs with intermediate age ($\approx 7-9$ Gyr) are found by Garro et al. (2022b) and Garro et al. (2022a), indicating that these objects are either real members of the bulge or intruders originating from the halo (Forbes & Bridges 2010; Massari et al. 2019; Romero-Colmenares et al. 2021). Moreover, spectroscopy of microlensed stars by Bensby et al. (2017) found a very wide age distribution for the bulge, with several star forming episodes at 3, 6, 8, and 11 Gyr. Also, they linked the ages to the metallicity of the microlensed stars. At [Fe/H] $\approx-0.5$ all stars appear to be older than 10 Gyr. At higher metallicities, the age ranges spans from the youngest one with $\sim 1$ Gyr to the oldest with $\sim 12-13$ Gyr. Additionally, the authors indicated that their age becomes younger with increasing metallicity. Below [Fe/H] $\approx-0.5$ essentially all stars are old, below solar but more metal-rich than $-0.5$ dex, the fraction is around 20%, and above solar metallicity more than one third of the stars are younger than 8 Gyr. Moreover, Bensby et al. (2017) found that Galactic bulge properties (MFD, star formation episodes, abundances trends) are similar to those of the MW thin and thick disks. However, it seems that the Bulge star formation rate was slightly faster than that of the local disk, as also proposed by Lucertini et al. (2022). Finally, their results again strengthen the observational evidence that supports the idea of a secular origin for the Galactic bulge, formed out of the other main stellar populations present in the central regions of our Galaxy.
@@
To summarise, it is really difficult to understand which of the three theories (hierarchical scenario, secular evolution or a combination of these two) best explains the formation and evolution of the bulge, as well as of the entire galaxy. Therefore, we prefer to use the GC population for trying to provide constrains, since these objects trace the Galaxy’s properties from the innermost regions towards the outskirts. In particular, we analysed numerous low-luminosity GCs, which probably survived strong dynamical processes, and we want to employ them in this paper in order to see what we can learn if we include this sample in the typical MW GC system.
2 The selected bulge Globular Cluster sample

### v2

# Over 200 globular clusters in the Milky Way and still none with super-Solar metallicity

@@ Abstract
Context. A large number of globular clusters in the Milky Way have been studied in recent years, especially in hidden regions such as those of the Galactic bulge.
Aims. The main goal of this work is to understand what we can learn if we include these new objects into the Milky Way globular cluster (GC) system that we know today. We compiled a catalog of 37 recently discovered globular clusters. Most of them are located in the Galactic bulge, but we also included some of the GCs for comparison.
Methods. We used a range of distributions for investigating the Galactic GC system based on the metallicity, luminosity function, and age. We considered three samples. We first treated the new GC sample separately from the known and well characterized GCs. Consequently, we merged these two samples, thereby upgrading the Milky Way GC system. Furthermore, we performed a comparison between our clusters sample and the field star population.
Results. We found a double-peaked distribution for the luminosity function, which shows an elongated faint end tail. Considering the ”merged” sample, the luminosity function peaks at $M_{V}^{up}=-7.00\pm 1.3$ mag and at $M_{V}^{up}=-4.1\pm 0.48$ mag. The metallicity distributions also display a bimodality trend. In this case, we compare our new sample compilation with previously published ones, finding that the distributions are in good general agreement. We also constructed the metallicity distribution for the field star sample and, by comparing it with that of the GCs, we learned that a high percentage of field stars show [Fe/H] $>0$; whereas we did not detect any GCs in the same metallicity range. To understand this inconsistency, we constructed the age-metallicity diagram for both samples, noting that the old and metal-poor population (age $\geq 8$ Gyr and [Fe/H] $\leq-1.0$) is represented by Gcs, while the young and metal-rich population (age $<8$ Gyr and [Fe/H] $>-1.0$) corresponds to field stars.
Conclusions. From the analysis of the GC luminosity function and metallicity distribution, we can conclude that many GCs, probably those that are very faint, have survived strong dynamical processes that are typical of the bulge regions. Moreover, we cannot exclude the possibility that some of them have been accreted during past merging events, especially the metal-poor component, whereas the metal-rich population may be related to the formation of the bulge and/or disk. Finally, the difference that we notice between the cluster and field star samples should be explored in the context of the evolutionary differences among these two stellar populations.
@@ 1 Introduction
In a cosmological context, the hierarchical galaxy formation process predicts that galaxies and their halos have a large number of progenitors at the formation redshift ($z\sim 3$) of globular clusters (GCs, De Lucia & Blaizot 2007; Lamers et al. 2017). With the advent of new observations from James Webb Space Telescope, new theories may be formulated and new observational discoveries can be established, which may change, confirm, or improve our knowledge. For instance, recent works have uncovered (mini-)quenched galaxies at redshifts of $z=4.4$ (Looser et al. 2023a) and $z=7.3$ (Looser et al. 2023b); thus, we cannot exclude the fact that GCs have been formed as well at these redshifts. Also, Mowla et al. (2022) found compact red companions to the Sparkler galaxy, which may be the first evolved GCs seen at $z_{spec}\approx 1.4$. They infer formation redshift of $z_{form}\sim 7-11$ for these candidates, corresponding to ages of $\sim 3.9-4.1$ Gyr at the epoch of observation and a formation time just $\sim 0.5$ Gyr after the Big Bang. Therefore, as these authors reported, if these ages are confirmed, at least some GCs appear to have formed contemporaneously with the large-scale reionization of the intergalactic medium, hinting at a deep connection between GC formation and the initial phases of galaxy assembly. Nevertheless, there is a long way to go before we can at least confirm the existence of those primordial GCs and begin constructing new models that describe these compact objects. This is especially true because their exact formation redshifts are somewhat uncertain for these unresolved and red objects, with no spectroscopic data yet available to offer clues to the real phase of these objects.
@@
Both simulations (e.g., Tonini 2013; Kruijssen 2015) and observations (e.g., Horta et al. 2023) have shown that much of the metal-poor GC population in massive galaxies was mostly accreted from cannibalized dwarf galaxies, whereas the metal-rich population formed preferentially in situ [^1] We use ”in situ” to refer to the GCs that formed during the initial collapse of the central galaxy, and ”ex-situ” for those one that are accreted during the subsequently mergers. or also likely during merging events (Hughes et al. 2019a; Garro et al. 2021a; Hasselquist et al. 2021). In this context, debris from accreted GCs and dwarf galaxies is expected to populate halo regions as well as the central regions of the MW (Das et al. 2020; Fernández-Trincado et al. 2019, 2021a). However, the number of confirmed merging footprints in the inner regions is still smaller than those found in the Galactic halo. For that reason, we want to investigate the Galactic bulge to collect clues to its formation and evolution history.
@@
Actually, there are many queries that remain unanswered regarding the bulge. For instance, it is not completely clear if the bulge is a simple classical bulge, formed from a monolithic collapse, and then grown up hierarchically by merging of galaxies (Guedes et al. 2013), or it was formed by secular evolution (Kormendy & Kennicutt 2004), so it is part of the peanut structure corresponding to the old Galactic thick disk. This would imply that the MW is a pure thin+thick disk galaxy, with only a possible limited contribution by a classical bulge (Di Matteo et al. 2015). Such a mix would represent a compound bulge (Athanassoula 2005; Erwin et al. 2014; Zoccali et al. 2014). Secondly, there are also queries about its evolution, such as how many accretions occurred, how the star formation history and so the chemical evolution proceeded, and so on.
@@
The main aim of this paper is to take advantage from bulge GC populations, which can be used as archaeological probes to reconstruct the formation and evolution of the MW. However, such regions as the Galactic bulge are very challenging to study, because of the high stellar crowding and differential reddening, which make the identification of new GCs a complicated task. Wide-field near-infrared (NIR) photometry such as the VISTA Variables in the Vía Láctea (VVV; Minniti et al. 2010; Saito et al. 2012) survey and its extension (VVVX; Minniti 2018) offers a huge advantage to detect several sparse and faint GCs heavily obscured by interstellar dust in the bulge region. Indeed, numerous ($\sim 300$) GC candidates have recently been discovered in these zones (Minniti et al. 2017, 2021c, 2021a; Dias et al. 2022; Garro et al. 2021b, 2022c). It seems that the number of known MW GCs (in total $N_{GC}\sim 170$ Vasiliev & Baumgardt 2021), but in particular the MW bulge GCs ($N_{GC}\sim 40$ Barbuy et al. 2018b, or maybe more as suggested by Bica et al. 2016) is too small if we take into account the mass of the MW and if we compare that number with similar systems, such as M31. In this context, for example, Abdullah & Kroupa (2018) (see their Figure 6) showed that the MW and M31 specific frequencies should be similar in their range of masses ($M/M_{\odot}=10^{10}-10^{11}$). Therefore, this may imply that the census of the MW clusters is still incomplete (Minniti et al. 2017). Indeed, searching and analysing new GCs in the MW bulge is tricky. They appear as round concentrations of stars, showing up as overdensities above the background in the optical and NIR images, but sometimes they are not real clusters but asterisms, background fluctuations, or possible streams of previously disrupted GCs. At the distance of the Galactic bulge, their typical half-light radii would range between 2 pc ($800^{\prime}.$80) for the smallest ones and 10 pc ($204^{\prime}.$20) for the largest ones. Additionally, since many of them are very faint and small objects, they are experiencing destruction processes, due to the fact that the deep potential well of the Galactic bulge strengthens the efficiency of such dynamic processes (Habibi et al. 2014; Khoperskov et al. 2018). Hence, this is why the Galactic bulge is referred to as the elephant graveyard (Minniti et al. 2017).
@@
Hence, we want to employ both known and well characterized GCs and the recently discovered ones for investigating the inner regions, using different tools: the metallicity distribution function (MD), the luminosity function (LF), age distributions, and the kinematics information. Additionally, we prefer expanding this analysis to the bulge field star (FS) populations, since it is expected that a small fraction of bulge FSs (proposed to be $\sim 5-10\%$Kruijssen 2012; Adamo et al. 2015) is born in bound systems, and then it is lost due to both merging events and/or destruction/dissolution processes (stellar evolution, two-body relaxation, tidal evaporation, and tidal shocks – Kruijssen et al. 2011; Madrid et al. 2017). Consequently, a comparison between these two populations may help us to understand and constrain the formation and evolution of the inner Galaxy. Nevertheless, these two populations are usually treated separately.
@@
With respect to what has been done thus far, different studies have reported that the Galactic bulge shows multiple peaks in its MD, features that are usually associated with different stellar populations (Ness et al. 2013; Bensby et al. 2017; Rojas-Arriagada et al. 2020), coexisting in the innermost regions. The existence of a metallicity gradient is demonstrated adopting different tracers (Minniti et al. 1995). Gonzalez et al. (2013) using the VVV photometry confirmed that metal-rich stars ([Fe/H] $\sim 0$) dominated the inner bulge in regions closer to the Galactic plane ($|b|<5$), while at larger scale heights, the mean metallicity of the bulge red giant population becomes significantly more metal-poor. Ness et al. (2013) and Debattista et al. (2017) found that the metal-rich population is dominant at low latitudes, but drops off rapidly with increasingly latitude. The metal-poor population instead follows the opposite trend, increasing toward high Galactic latitudes. The intermediate-metallicity population dominates everywhere except near the disk, where the metal-rich population dominates. On the other hand, works by Barbuy et al. (2018a) and Savino et al. (2020) found a predominance of moderately metal-poor populations in the innermost Galactocentric regions, which extends to large distances above the Galactic plane. This is probably related to the fact that the old and intermediate age RR Lyrae populations were taken into account and they are not expected to exhibit a strong peanut shape, as red clump stars (Gonzalez et al. 2015), since they should be older than the formation of the bar itself (Debattista et al. 2017). On the other side, if we take into account GCs, we find the metal-rich ones especially in the inner regions of the Galaxy, while the metal-poor one in the halo.
@@
We may note that the meanings of ”metal-poor” and ”metal-rich” are not strictly defined and depend on the different contexts where these terms are used. For instance, for the ”metal-rich” FSs we often refer to solar metallicity ([Fe/H] $\approx 0$), while for the ”metal-rich” GCs, we consider values of [Fe/H] $>-0.5$, which might be referred to as ”intermediate-metallicity” for the field overall. Therefore, in the present work, we broadly define ”metal-poor” as [Fe/H] $<-1.0$, whereas ”metal-rich” if [Fe/H] $\geq-1.0$, both for FSs and GCs. Moreover, we refer to ”super metal-rich” or ”super-Solar” to indicate those populations with [Fe/H] $\geq 0$. Various nomenclatures are often used in the literature, such as ”moderately metal-poor” (which may correspond to $-2.0\lesssim$ [Fe/H] $\lesssim-1.0$) or ”intermediate metallicity” (which may be equivalent to $-1.5\lesssim$ [Fe/H] $\lesssim-0.6$); however, this depends once again on the populations involved. We prefer not to use this nomenclature.
@@
Today, we know of $\sim 170$ well-characterized GCs in the MW (Vasiliev & Baumgardt 2021). Broadly, the Galactic GC system shows a bimodal metallicity distribution, which indicates that in the MW coexist two major subpopulations, but always the metallicity is typically sub-Solar ([Fe/H] $<0$). In spite of that, there are some GCs that may show [Fe/H] $>0$. This is not completely demonstrated, and the misgivings are strongly related to the abundance scale’s calibrations adopted. Therefore, we cannot ignore them. For example, Carretta et al. (2009) suggested a [Fe/H] $>+0.40\pm 0.09$ for Liller 1 GC. Recently, Dalessandro et al. (2022) and Crociati et al. (2023) showed a bimodal iron distribution, with a sub-solar component ([Fe/H] $=-0.48$ dex with $\sigma=0.22$, and age $\approx 12$ Gyr) and a super-solar component ([Fe/H] $=+0.26$ dex with $\sigma=0.17$, and age $\approx 1-3$ Gyr) for Liller 1. For this purpose, they suggested that Liller 1 experienced three main bursts and a low, but constant, activity of star formation over the entire lifetime. Also, NGC 6528 may have a super-solar metallicity as found by Carretta et al. (2009) ([Fe/H] $=+0.07\pm 0.08$) and Liu et al. (2017) ([Fe/H] $=+0.04\pm 0.07$), but in both cases, the uncertainties are greater than the value itself; therefore, we cannot definitively conclude either way whether we indeed have a case of a super metal-rich GC. It is clear that the analysis of the bulge MD is very complex, since this strongly depends on target selection, tracers, and analysis methods.
@@
Moreover, the $\alpha$-elements are efficient tracers of Galactic evolution, especially when [$\alpha$/Fe]-[Fe/H] diagrams are used. An $\alpha$-enhanced plateau at low [Fe/H] is typical both for the MW and external galaxies, this is due to the fast enrichment from massive ($408<M/M_{\odot}<$ 40) stars that explode as type II supernovae (SNII), while a following decline is expected once low-mass (in binary systems with $81.4<M/M_{\odot}<$ 8) stars begin to explode as type Ia supernovae (SNIa). This creates the ”knee” point, which can move to higher metallicity for higher star formation rates (Matteucci & Greggio 1986). For example, Queiroz et al. (2021) found an apparent chemical bimodality in key abundance ratios, such as [$\alpha$/Fe], [C/N], and [Mn/O], which probe different enrichment timescales, indicating a star formation gap between the $\alpha$-rich and $\alpha$-poor populations. They also found signatures of the bar, a thin inner disk, a thick inner disk, and a broad metallicity population with large velocity dispersion indicative of a pressure-supported component. Furthermore, they found a group of counter-rotating stars that could be an accreted metal-poor population originated during a gas-rich accretion phase in the early formation of the bulge, or of a clumpy star formation during the earliest phases of the early disk that migrated into the bulge (Fiteni et al. 2021). In summary, many observations suggest that the formation of the boxy bulge occurred later, originating from the vertical instability of the Galactic bulge, formed via secular evolution from the stellar disk. In this scenario, the old and metal-poor stellar population dominates at higher scale heights due to the non-mixed orbits of originally hotter thick disk stars. Lian et al. (2020) derived three-phase star formation history of the MW bulge, which consists of an initial starburst, followed by a rapid star formation quenching episode and a lengthy, quiescent secular evolution phase. In fact, they found that the metal-poor, high-$\alpha$ bulge stars ([Fe/H] $<0.0$ and [Mg/Fe] $>0.15$) were formed rapidly ($<2$ Gyr) during the early starburst. The density gap between the high- and low-$\alpha$ sequences is due to the quenching process. The metal-rich, low-$\alpha$ population ([Fe/H] $>0.0$ and [Mg/Fe] $<0.15$) then accumulates gradually through inefficient star formation during the secular phase. This is qualitatively consistent with the early star formation history (SFH) of the inner disk. Additionally, this explains why there is a fraction of young stars with age $<5$ Gyr in the bulge. Indeed, we know that the bulge mean age is about 10 Gyr, with a lower-limit of $\sim 5$ Gyr (Clarkson et al. 2011), where well-mixed clusters and stellar populations coexist. The typical age of GCs in the innermost regions is older than 10 Gyr (Harris 1996; Minniti 1995), however GCs with intermediate age ($\approx 7-9$ Gyr) are found by Garro et al. (2022b) and Garro et al. (2022a), indicating that these objects are either real members of the bulge or intruders originating from the halo (Forbes & Bridges 2010; Massari et al. 2019; Romero-Colmenares et al. 2021). Moreover, spectroscopy of microlensed stars by Bensby et al. (2017) found a very wide age distribution for the bulge, with several star forming episodes at 3, 6, 8, and 11 Gyr. Also, they linked the ages to the metallicity of the microlensed stars. At [Fe/H] $\approx-0.5$ all stars appear to be older than 10 Gyr. At higher metallicities, the age ranges spans from the youngest one with $\sim 1$ Gyr to the oldest with $\sim 12-13$ Gyr. Additionally, the authors indicated that their age becomes younger with increasing metallicity. Below [Fe/H] $\approx-0.5$ essentially all stars are old, below solar but more metal-rich than $-0.5$ dex, the fraction is around 20%, and above solar metallicity more than one third of the stars are younger than 8 Gyr. Moreover, Bensby et al. (2017) found that Galactic bulge properties (MFD, star formation episodes, abundances trends) are similar to those of the MW thin and thick disks. However, it seems that the bulge star formation rate was slightly faster than that of the local disk, as also proposed by Lucertini et al. (2022). Finally, their results again strengthen the observational evidence that supports the idea of a secular origin for the Galactic bulge, formed out of the other main stellar populations present in the central regions of our Galaxy.
@@
To summarize, it really is difficult to understand which of the three theories (hierarchical scenario, secular evolution, or a combination of these two) best explains the formation and evolution of the bulge, as well as that of the entire galaxy. Therefore, we prefer to use the GC population in our attempts to provide constrains, since these objects trace the Galaxy’s properties from the innermost regions toward the outskirts. In particular, we analyzed numerous low-luminosity GCs, which probably survived strong dynamical processes. We employ them in this paper to see what we can learn if we include this sample in the typical MW GC system.
2 Sample of selected bulge GCs